OpenStudy (deadshot):

Simplify quantity 12 x minus 96 over quantity x squared minus 4x minus 32 all over quantity 6 x squared minus 6 x minus 12 over 3 x plus 3 ?

OpenStudy (deadshot):

\[\frac{ \frac{ 12x-96 }{ x^2-4x-32 } }{ \frac{ 6x^2-6x-12 }{ 3x+3 } }\]

OpenStudy (anonymous):

\[\frac{ 12x-96 }{ x^2-4x-32 }*\frac{ 3x+3 }{ 6x^2-6x-13 }\]

OpenStudy (anonymous):

can you solve it now

OpenStudy (deadshot):

so, it's \[\frac{ 36x^2-288 }{ 6x^4+24x^2+416 }\] right?

OpenStudy (anonymous):

no dont solve it like that

OpenStudy (anonymous):

factorize the terms

OpenStudy (deadshot):

So then it's \[\frac{ 18x^2-144 }{ 3x^4+12x^2+208 }\]

OpenStudy (anonymous):

oh god do you know factorization?

OpenStudy (deadshot):

Yes, you find the gcf and remove it from the terms, right?

OpenStudy (anonymous):

-_-"

OpenStudy (anonymous):

12x-96 how else can you write this?

OpenStudy (deadshot):

12(x-8)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

and 3x+3?

OpenStudy (deadshot):

3(x+1)

OpenStudy (anonymous):

now \[x^2-4x-32\] can you factorize this?

OpenStudy (deadshot):

(x-8)(x+4)

OpenStudy (anonymous):

very good and the last term \[6x^2-6x-13\]

OpenStudy (deadshot):

6(x+1)(x-2)

OpenStudy (anonymous):

it was 12 not 13 sorry and good now put these values in the equation i gave you and solve it what you get?

OpenStudy (deadshot):

\[\frac{ 12(x-8) }{ (x-8)(x+4) }\times \frac{ 3(x+1) }{ 6(x+1)(x-2) }\]

OpenStudy (anonymous):

now cancel the terms x+1 ,x-8 gets canceled and 2 times 6 is 12 so 6 gets canceled and 12 becomes 2 tell me the final answer

OpenStudy (deadshot):

\[\frac{ 6 }{ x^2+2x-8 }\]

OpenStudy (anonymous):

dont solve the brackets \[\frac{ 6 }{ (x+4)(x-2) }\] is your answer

OpenStudy (deadshot):

Oh, okay, Thanks!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

and p.s. \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (deadshot):

Okay, Thanks!

OpenStudy (anonymous):

:)